OpenStudy (anonymous):

Icebergs consist of freshwater ice and float in the ocean with only about 10.0 of their volume above water (the "tip of the iceberg," so to speak). This percentage can vary, depending on the condition of the ice. Assume that the ice has the density given in Table 13.1, although, in reality, this can vary considerably, depending on the condition of the ice and the amount of impurities in it. What does this 10.0 tell us is the density of seawater? What percentage of the icebergs' volume would be above water if they were floating in a large freshwater lake such as Lake Superior?

OpenStudy (anonymous):

it tells you that seawater is 10% denser and in freshwater ice would float but it wouldnt stick out, again depending on the ice and the temperature and so forth...

OpenStudy (anonymous):

Use the Force rapresentation, the iceberg had gravity applied to it's center of mass, and the archimedes force applied to the part of iceberg underwater: mg=m*R*V Fa=m*r*v where R are the density of ice, r the density of water, V the volume of the ice, v the volume of the ice underwater, so we have: mRV-mrv=0 (equilibrium) mRV=mrv => v/V=R/r that is the percentile we want, so just divide the density of ice from the density of water, and et volià ;) nice to help you!